# Harmony: A Joint Self-Supervised and Weakly-Supervised Framework for Learning General Purpose Visual Representations

### v1

@@ 3.2 Feature Self-distillation
On top of the contrastive objective, $\mathcal{L}_{Contrastive}$, we add a self-supervised feature self-distillation loss following iBOT [55], which we find can significantly boost performance across all downstream tasks evaluated (see Table 4). This loss consists of both global and local objectives that go hand-in-hand to learn generalized visual features. Conceptually, the goal of this loss in our framework is to learn visual features that might not be described in the caption of the contrastive loss, and learn more localized features by adding a local objective that operates at the patch level.
Global objective. Following [6, 38, 20, 55] we utilize a teacher encoder $\bar{E}_{V}$ that defined as being the EMA a student encoder ${E}_{V}$, which is trained with gradient optimization. The teacher and student models, $\bar{E}_{V}$ and ${E}_{V}$, are the same models used in the contrastive loss from Section 3.1. Just like in [6, 55], we attach a multi-layer perceptron (MLP) projection head $h$, such that $h:g=h\circ f$, where $f$ is either $\bar{E}_{V}$ or ${E}_{V}$, resulting in $\bar{g}$ and $g$ models, respectively. Given an input image $x$ we generate two different augmentations, $x_{1}$ and $x_{2}$, defined as random crops or views of $x$. We feed these two augmentations to $\bar{g}$ and $g$, and optimize the following loss function:
$\mathcal{L}_{\text{CLS}}=-\frac{1}{2}\left[\mathcal{H}(\bar{P}(x_{1}),P(x_{2}))+\mathcal{H}(\bar{P}(x_{2}),P(x_{1})\right],$ (6)
@@
Local objective. Equation 6 is applied on the class token (CLS) of the vision transformer [18], which is why we refer to it as $\mathcal{L}_{\text{CLS}}$. Since CLS tokens aggregate information across different patches into a single token, minimizing $\mathcal{L}_{\text{CLS}}$ learns more global features, sometimes disregarding granular details that are helpful for dense prediction tasks like semantic segmentation. On top of this global-level objective we aim to learn localized features by employing a patch-level loss using masked image modeling (MIM), following iBOT [55]. More precisely, given the sequence of image tokens $x=\{\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\mathrm{x}_{N}\}$ being processed by a ViT, we sample from the masking set $m\in\{0,1\}^{N}$ according to a ratio $r$. If $\mathrm{m_{i}}=1$, we replace the original $\mathrm{x_{i}}$ with a special mask token $\mathrm{x}_{m}$, resulting in a masked view $\hat{x}$ formalized as $\hat{x}\triangleq\left\{\mathrm{\hat{x}}_{i}\left((1-\mathrm{m}_{i})\mathrm{x}_{i}+\mathrm{m}_{i}\mathrm{x}_{m}\right)\right\}_{i=1}^{N}$ [55]. We subsequently feed $\hat{x}$ and $x$ to $\bar{g}$ and $g$, respectively, optimizing the loss function over all patch tokens:
$\mathcal{L}_{\text{MIM}}=-\sum_{i=1}^{N}\,\mathrm{m_{i}}\,\mathcal{H}(\bar{P}(\hat{\text{x}}_{i}),P(\text{x}_{i}))\,.$ (8)
As shown in [55], adding the patch-level loss $\mathcal{L}_{\text{MIM}}$ on top of $\mathcal{L}_{\text{CLS}}$, which is originally from DINO [6], improves downstream performance on both dense prediction and classification tasks.
Moreover, we also minimize the cross entropy loss between the embeddings of $x_{1}$ and $x_{2}$ extracted by the teacher model $\bar{E}_{V}$ and embeddings of smaller, local crops $\{y_{1},y_{2},\dots,y_{L}\}$ extracted by the student model ${E}_{V}$ and generated using a multi-crop augmentation strategy [5], which is shown to increase performance in [6, 55]. This multi-crop optimization is added only to $\mathcal{L}_{\text{CLS}}$.
The final self-distillation loss $\mathcal{L}_{distillation}$ is as the average of $\mathcal{L}_{\text{CLS}}$ and $\mathcal{L}_{\text{MIM}}$, i.e.,
$\mathcal{L}_{Distillation}=\frac{\mathcal{L}_{\text{CLS}}+\mathcal{L}_{\text{MIM}}}{2}\,.$ (9)
@@ 3.3 Pixel Reconstruction
The $\mathcal{L}_{Distillation}$ loss in Section 3.2 is applied on the feature space. On top of this, we add another pixel level loss for learning more granular features, which we find slightly improve performance for segmentation tasks (see Table 4). Specifically, we follow the approach in MAE [24] in that, given a sequence of patch token $\{\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\mathrm{x}_{N}\}$, we mask (remove) $P$ number of the patch tokens at random (as opposed to replacing them with $x_{m}$ like in Section 3.2). We feed the remaining $(L-P)+1$ of tokens, where the $1$ is the $x_{\text{CLS}}$ token and $L$ is the total number of patches in the original images, to ${E}_{V}$, which is the same encoder used in the contrastive objective (Section 3.1) and feature self-distillation (Section 3.2). We end up with a sequence of image embeddings $e=\{\mathrm{e}_{\text{CLS}},\mathrm{e}_{1},\mathrm{e}_{2},\ldots,\mathrm{e}_{M}\}$, where $M=L-P+1$. We then add $P$ number of mask tokens $\text{x}_{m}$ to $e$ in the same position (or index) they were removed from to obtain the masked embedding $\hat{e}$. Subsequently, $\hat{e}$ is passed to a decoder $D_{V}$, which will up-sample the embeddings from $d_{D}$ to the $H\times W$, where $d_{D}$ is the embedding size of $D_{V}$, $H$ and $W$ are the height and width of the original patch size, respectively. We set $P=L\times 0.75$, following [24].
The loss is then calculated as the Mean Squared Error between the predicted pixels, $p_{i}$, and the L2 normalized pixels of the original target in a patch, $\hat{p_{i}}$. In other words, the loss becomes:
$\mathcal{L}_{Reconstruction}=\frac{1}{L}\sum_{i=1}^{L}\,\mathrm{m_{i}}\,(p_{i}-\hat{p}_{i})^{2}\,.$ (10)
@@ 3.4 MLM and Text Self-distillation
Even though the goal of this work is visual representation learning, we hypothesize that improving the textual representations in our text encoder ${E}_{T}$ can improve visual representation in the vision encoder ${E}_{V}$ both by making the soft learning targets described in Section 3.1 more meaningful, and by improving the overall contrastive learning task. Moreover, improving textual representations can have a significant impact on zero-shot evaluations as supported by results from [17] and our Table 4. As a result, we added two textual objectives that the text encoder ${E}_{T}$ will now optimize on top of the contrastive loss. These objectives are masked language modeling and text self-distillation.
Masked language modeling. The first additional textual objective is masked language modeling, as described in BERT [15], which predicts masked words given the context of surrounding words. MLM can be viewed analogously as pixel prediction, in the sense that both optimize outputs that are in the same level as original inputs (tokenized words in the context of MLM and pixels in pixel prediction).
We feed a masked view $\hat{t}$ of a tokenized caption $t$ to the student text encoder ${E}_{T}$, where word tokens are randomly masked using a Bernoulli distribution $m\sim\text{Bernoulli}(p)$ with $p=0.2$, with $m$ being the mask vector that is applied on a caption $t$ to generate $\hat{t}$. Unlike BERT [15], we don’t replace words with random words, nor keep some predicted words unchanged. Masked word tokens are replaced with a mask token $t_{\text{m}}$ to generate $\hat{t}$. Our final MLM loss is then formulated as
$\mathcal{L}_{MLM}=\sum_{i=1}^{C}\,\mathrm{m_{i}}\,\mathcal{H}(D_{T}({E}_{T}(\hat{t}_{i})),t_{i})\,,$ (11)
@@
Text self-distillation. On top of the MLM loss, we add a self-distillation loss by utilizing the student and teacher text encoders, ${E}_{T}$ and $\bar{E}_{T}$, motivated by iBOT’s patch-level objective [55]. Unlike the MLM loss, this objective will function at the embedding level, which offers softer targets for the student text encoder compared to MLM. We attach MLP projection heads to ${E}_{T}$ and $\bar{E}_{T}$, resulting in $g$ and $\bar{g}$, respectively. We then optimize the equation
$\mathcal{L}_{TextDistillation}=-\sum_{i=1}^{C}\,\mathrm{m_{i}}\,\mathcal{H}(\bar{P}(\hat{t_{i}}),P(t_{i}))\,,$ (12)
@@ 3.5 Harmony’s Objective
Our finalized framework, Harmony, optimizes the five described objectives simultaneously. In other words, our final loss is a linear transformation of all five losses or the equation:
$\mathcal{L}_{H}=\mathcal{L}_{Contrastive}+\alpha\mathcal{L}_{Distillation}+\beta\mathcal{L}_{Reconstruction}+\gamma\mathcal{L}_{MLM}+\delta\mathcal{L}_{TextDistillation}\,.$ (13)
@@ 4 Experiments
Here, we experimentally evaluate Harmony against the baseline methods CLIP [28], iBOT [55], and MAE [24], as well as previously leading joint SSL and WSL methods, MaskCLIP [17] and SLIP [37]. We start in Section 4.1 by describing our model architecture and training setup, then we present our results in 4.2, and finally end with an ablation and hyper-parameter tuning study in 4.3.
@@ 4.1 Training Setup
Model architecture. Six transformer networks [46], three of which are vision transformers [18], of different sizes and configurations, make up our final framework: (1, 2) student and teacher vision encoders, ${E}_{V}$ and $\bar{E}_{V}$; (3, 4) student and teacher text encoders, ${E}_{T}$ and $\bar{E}_{T}$; (5) vision decoder $D_{V}$; and (6) text decoder $D_{T}$. For the vision encoders, ${E}_{V}$ and $\bar{E}_{V}$, we use the ViT-S/16, which consists of 12 layers, 384 embedding dimensions, and 6 attention heads. For the text encoders, ${E}_{T}$ and $\bar{E}_{T}$, we use 12 layers, 512 embedding dimensions, and 8 heads, following CLIP [40]. The number of text tokens is fixed to 77 with necessary truncations or paddings. Moreover, following [24], we define our vision decoder $D_{V}$ as a 8-layer, 512-embedding dimension, and 16-head ViT. Finally, our text decoder $D_{T}$, is made of up 4 layers, 512 embedding dimension, and 8 heads.
Pre-training. We pre-train a ViT-S/16 using Harmony on CC3M for 25 epochs. We use a batch size of 768, and an AdamW optimizer [33] with an $\epsilon$ value of $1e^{-6}$ to improve training stability using mixed precision [35]. We set the number of global crops per a given image to 2 and local crops to 8. We use a masking ratio of 75% for the pixel prediction task and pre-train on up but to 16 32-GB V100 cards for most trainings. Additionally, we pre-train a ViT-B/16 using the same settings, but with 6 local crops and a batch size of 912, using 24 V100s, to show the scaling capabilities of Harmony. Further information about our pre-training setup can be found in Appendix A.
Downstream. To asses the generalizability of learned features using our proposed method, we evaluate our model on a variety of datasets, including ImageNet-1k [41] for fine-tuning, linear-probing, and zero-shot classifications, ADE20K [54] for semantic segmentation, MS-COCO [32] for object detection and instance segmentation, Flicr30K [52] and MS-COCO for image-text retrieval, as well as many other datasets for zero-shot evaluations. We fine-tune for 100 epochs on ImageNet-1k and for 160k iterations for ADE20K, using UperNet [49]. For object detection and instance segmentation, we use a Cascade Mask R-CNN [4] and fine-tune for 12 epochs. Further details on downstream settings can be found in Appendix A.
@@ 4.2 Results
Table 1: Comparing Harmony against baseline SSL and previous best methods. All methods are pre-trained on CC3M for 25 epochs. We used the same pre-training settings for MAE iBOT, CLIP, and Harmony. To reproduce SLIP, we used their public code base [37]. We reproduced MaskCLIP ourselves, since their code was not public (see appendix C for further detail).
Method Size INET-1K (ACC) ADE20K (mIoU) MS-COCO (mAP)
ViT 0-Shot FT LIN SSEG DET ISEG
MAE [24] S 70.9 14.6 26.9 28.9 25.6
iBOT [55] S 79.5 58.2 37.0 44.9 39.0
CLIP [40] S 16.7 76.8 56.4 35.3 40.7 35.5
SLIP [37] S 18.1 76.5 51.5 37.8 40.0 35.0
MaskCLIP [17] S 17.7 77.0 56.1 34.7 39.4 34.5
Harmony S 21.0 79.8 64.6 37.9 45.8 39.8
Harmony B 23.4 82.3 69.8 45.5 48.3 41.7
@@
We begin our evaluation of Harmony by analyzing its downstream performance for a variety of visual tasks under different settings. We start in Table 1 by presenting our main result for classification, segmentation, and detection tasks. We show that Harmony significantly outperforms CLIP, and outperforms iBOT [55] and MAE [24] across all high-level and low-level tasks evaluated. We also show that Harmony outperforms both SLIP [37] and MaskCLIP [17], which are joint methods that are similar to ours. Interestingly, MAE performs substantially worse than the other methods, which could be due to the distribution of the web-scraped data in CC3M. Even then, using the pixel reconstruction objective in MAE still compliments features for Harmony (shown as an ablation in Table 4). Moreover, SLIP performs worse than expected on classification tasks, which could be due to the small batch size for SimCLR’s contrastive loss [9] (768 vs 4096 in the original SLIP [37]). This issue is less prominent in methods that use self-distillation like Harmony and MaskCLIP [17] because they do not rely on negative examples in the batch.
@@
Zero-shot evaluations across datasets. Moreover, we compare the zero-shot capabilities of Harmony to CLIP, SLIP, and MaskCLIP across datasets of different sizes. We focus on general and natural datasets (as opposed to domain specific like StanfordCars [50] and FGVC-Aircraft [34]), and ImageNet robustness datasets like ImageNet-R. We present our results in Table 2. Harmony outperforms the other methods in 7 out of the 9 datasets, sometimes by relatively large margins – around +10% on Caltech-101 and +5% on Kinetics. We show qualitative examples in Appendix E.
@@
Retrieval. We also evaluate Harmony for image-text zero-shot retrievals on MS-COCO [32] and Flickr30K [52]. We observe that for top 5 (R@5) and top 10 retrievals, the gap between Harmony and the other methods grows rapidly, indicating that our method is more robust, since it still generates higher probabilities for the correct retrieval, even if it can’t retrieve it within the first examples.
@@
Table 2: Zero-shot evaluations. We compare Harmony against CLIP, SLIP, and MaskCLIP on general and natural image benchmarks for 0-shot classification. Harmony outperforms the other methods in 7 out of 9 datasets chosen.
Size
Average
Caltech-101 [30]
STL-10 [13]
Kinetics700 [7]
MNIST [29]
CLEVR [28]
ImageNet-A [26]
ImageNet-O [26]
ImageNet-R [25]
ImageNet [41]
CLIP [40] S 26.2 49.6 86.5 15.4 9.8 12.3 3.9 23.1 18.2 16.7
SLIP [37] S 27.4 48.9 86.0 15.3 13.1 15.7 4.7 24.7 20.2 18.1
MaskCLIP [17] S 25.6 47.8 82.3 14.6 9.7 12.1 3.9 23.7 18.4 17.7
Harmony S 30.0 57.7 92.6 20.3 9.8 12.5 5.8 25.0 25.7 21.0
Harmony B 32.2 61.0 94.0 22.5 11.7 12.6 7.8 25.9 30.9 23.4

### v3

@@ 3.2 Feature Self-distillation
On top of the contrastive objective $\mathcal{L}_{C}$, we add a self-supervised feature self-distillation loss following iBOT (Zhou et al., 2022), which we find can significantly boost performance across all downstream tasks evaluated (see Table 4). This loss consists of both global and local objectives that go hand-in-hand to learn generalized visual features. Conceptually, the goal of this loss in our framework is to learn visual features that might not be described in the caption of the contrastive loss, and learn more localized features by adding a local objective that operates at the patch level.
Global objective. Following (Caron et al., 2021b; Oquab et al., 2024; Grill et al., 2020; Zhou et al., 2022) we utilize a teacher encoder $\bar{E}_{V}$ that is defined as the EMA of a student encoder ${E}_{V}$, which is trained with gradient optimization. The teacher and student models, $\bar{E}_{V}$ and ${E}_{V}$, are the same models used in the contrastive loss from Section 3.1. Just like in (Caron et al., 2021b; Zhou et al., 2022), we attach a multi-layer perceptron (MLP) projection head $h$, such that $h:g=h\circ f$, where $f$ is either $\bar{E}_{V}$ or ${E}_{V}$, resulting in $\bar{g}$ and $g$ models, respectively. Given an input image $x$ we generate two different augmentations, $x_{1}$ and $x_{2}$, defined as random crops or views of $x$. We feed these two augmentations to $\bar{g}$ and $g$, and optimize the following loss function:
$\mathcal{L}_{\text{CLS}}=-\frac{\mathcal{H}(\bar{P}(x_{1}),P(x_{2}))+\mathcal{H}(\bar{P}(x_{2}),P(x_{1}))}{2},$ (6)
@@
Local objective. Equation 6 is applied on the class token (CLS) of the vision transformer (Dosovitskiy et al., 2021), which is why we refer to it as $\mathcal{L}_{\text{CLS}}$. Since CLS tokens aggregate information across different patches into a single token, minimizing $\mathcal{L}_{\text{CLS}}$ learns more global features, sometimes disregarding granular details that are helpful for dense prediction tasks like semantic segmentation. On top of this global-level objective we aim to learn localized features by employing a patch-level loss using masked image modeling (MIM), following iBOT (Zhou et al., 2022). More precisely, given the sequence of image tokens $x=\{\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\mathrm{x}_{N}\}$ being processed by a ViT, we sample from the masking set $m\in\{0,1\}^{N}$ according to a ratio $r$. If $\mathrm{m_{i}}=1$, we replace the original $\mathrm{x_{i}}$ with a special mask token $\mathrm{x}_{m}$, resulting in a masked view $\hat{x}$ formalized as $\hat{x}\triangleq\left\{\mathrm{\hat{x}}_{i}\left((1-\mathrm{m}_{i})\mathrm{x}_{i}+\mathrm{m}_{i}\mathrm{x}_{m}\right)\right\}_{i=1}^{N}$ (Zhou et al., 2022). We subsequently feed $x$ and $\hat{x}$ to $\bar{g}$ and $g$, respectively, optimizing the loss function over all patch tokens:
$\mathcal{L}_{\text{MIM}}=-\sum_{i=1}^{N}\,\mathrm{m_{i}}\,\mathcal{H}(\bar{P}(\text{x}_{i}),P(\hat{\text{x}_{i}}))\,.$ (8)
As shown in (Zhou et al., 2022), adding the patch-level loss $\mathcal{L}_{\text{MIM}}$ on top of $\mathcal{L}_{\text{CLS}}$, which is originally from DINO (Caron et al., 2021b), improves downstream performance on both dense prediction and classification tasks.
Moreover, we also minimize the cross entropy loss between the embeddings of $x_{1}$ and $x_{2}$ extracted by the teacher model $\bar{E}_{V}$ and embeddings of smaller, local crops $\{y_{1},y_{2},\dots,y_{L}\}$ extracted by the student model ${E}_{V}$ and generated using a multi-crop augmentation strategy (Caron et al., 2021a), which is shown to increase performance in (Caron et al., 2021b; Zhou et al., 2022). This multi-crop optimization is added only to $\mathcal{L}_{\text{CLS}}$.
The final self-distillation loss $\mathcal{L}_{D}$ is as the average of $\mathcal{L}_{\text{CLS}}$ and $\mathcal{L}_{\text{MIM}}$, i.e.,
$\mathcal{L}_{D}=\frac{1}{2}(\mathcal{L}_{\text{CLS}}+\mathcal{L}_{\text{MIM}})\,.$ (9)
@@ 3.3 Pixel Reconstruction
The $\mathcal{L}_{D}$ loss is applied on the feature space. On top of this, we add another pixel level loss for learning more granular features, which we find slightly improve performance for segmentation tasks (see Table 4). Specifically, we follow the approach in MAE (He et al., 2021) in that, given a sequence of patch token $\{\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\mathrm{x}_{N}\}$, we mask (remove) $P$ number of the patch tokens at random (as opposed to replacing them with $x_{m}$ like in Section 3.2). We feed the remaining $(L-P)+1$ of tokens, where the $1$ is the $x_{\text{CLS}}$ token and $L$ is the total number of patches in the original images, to ${E}_{V}$, which is the same encoder used in the contrastive objective (Section 3.1) and feature self-distillation (Section 3.2). We end up with a sequence of image embeddings $e=\{\mathrm{e}_{\text{CLS}},\mathrm{e}_{1},\mathrm{e}_{2},\ldots,\mathrm{e}_{M}\}$, where $M=L-P+1$. We then add $P$ number of mask tokens $\text{x}_{m}$ to $e$ in the same position (or index) they were removed from to obtain the masked embedding $\hat{e}$. Subsequently, $\hat{e}$ is passed to a decoder $D_{V}$, which will up-sample the embeddings from $d_{D}$ to the $H\times W$, where $d_{D}$ is the embedding size of $D_{V}$, $H$ and $W$ are the height and width of the original patch size, respectively. We set $P=L\times 0.75$, following He et al. (2021).
The loss is then calculated as the Mean Squared Error between the predicted pixels $p_{i}$ and the L2 normalized pixels of the original target in a patch $\hat{p_{i}}$. In other words, the loss becomes:
$\mathcal{L}_{R}=\frac{1}{L}\sum_{i=1}^{L}\,\mathrm{m_{i}}\,(p_{i}-\hat{p}_{i})^{2}\,.$ (10)
@@ 3.4 MLM and Text Self-distillation
Even though the goal of this work is visual representation learning, improving textual representations can have a significant impact on zero-shot evaluations as supported by results from Dong et al. (2023) and our Table 4. As a result, we added two textual objectives that the text encoder ${E}_{T}$ will now optimize on top of the contrastive loss. These objectives are masked language modeling and text self-distillation.
Masked language modeling. The first additional textual objective is masked language modeling, as described in BERT (Devlin et al., 2019), which predicts masked words given the context of surrounding words. MLM can be viewed analogously as pixel prediction, in the sense that both optimize outputs that are in the same level as original inputs (tokenized words in the context of MLM and pixels in pixel prediction).
We feed a masked view $\hat{t}$ of a tokenized caption $t$ to the student text encoder ${E}_{T}$, where word tokens are randomly masked using a Bernoulli distribution $m\sim\text{Bernoulli}(p)$ with $p=0.2$, with $m$ being the mask vector that is applied on a caption $t$ to generate $\hat{t}$. Unlike BERT (Devlin et al., 2019), we don’t replace words with random words, nor keep some predicted words unchanged. Masked word tokens are replaced with a mask token $t_{\text{m}}$ to generate $\hat{t}$. Our final MLM loss is then formulated as
$\mathcal{L}_{M}=\sum_{i=1}^{C}\,\mathrm{m_{i}}\,\mathcal{H}(t_{i},D_{T}({E}_{T}(\hat{t}_{i})))\,,$ (11)
@@
Text self-distillation. On top of the MLM loss, we add a self-distillation loss by utilizing the student and teacher text encoders, ${E}_{T}$ and $\bar{E}_{T}$, motivated by iBOT’s patch-level objective (Zhou et al., 2022). Unlike the MLM loss, this objective will function at the embedding level, which offers softer targets for the student text encoder compared to MLM. We attach MLP projection heads to ${E}_{T}$ and $\bar{E}_{T}$, resulting in $g$ and $\bar{g}$, respectively. We then optimize the equation
$\mathcal{L}_{TD}=-\sum_{i=1}^{C}\,\mathrm{m_{i}}\,\mathcal{H}(\bar{P}(t_{i}),P(\hat{t_{i}}))\,,$ (12)
@@ 3.5 Harmony’s Objective
Our finalized framework, Harmony, optimizes the five described objectives simultaneously. In other words, our final loss is a linear combination of all five losses or the equation:
$\mathcal{L}_{H}=\mathcal{L}_{C}+\alpha\mathcal{L}_{D}+\beta\mathcal{L}_{R}+\gamma\mathcal{L}_{M}+\delta\mathcal{L}_{TD}\,.$ (13)
@@ 4 Experiments
Here, we experimentally evaluate Harmony against the baseline methods CLIP (Johnson et al., 2016), SigLIP (Zhai et al., 2023), iBOT (Zhou et al., 2022), and MAE (He et al., 2021), as well as previously leading joint SSL and WSL methods, SLIP (Mu et al., 2021), SILC (Naeem et al., 2024), MaskCLIP (Dong et al., 2023), and DetailCLIP (Monsefi et al., 2024). We start in Section 4.1 by describing our model architecture and training setup, then we present our results in Section 4.2, and finally end with an ablation and hyper-parameter tuning study in Section 4.3.
@@ 4.1 Training Setup
Model architecture. Six transformer networks (Vaswani et al., 2023), three of which are vision transformers (Dosovitskiy et al., 2021), of different sizes and configurations, make up our final framework: (1, 2) student and teacher vision encoders, ${E}_{V}$ and $\bar{E}_{V}$; (3, 4) student and teacher text encoders, ${E}_{T}$ and $\bar{E}_{T}$; (5) vision decoder $D_{V}$; and (6) text decoder $D_{T}$. For the vision encoders, ${E}_{V}$ and $\bar{E}_{V}$, we use ViT sizes small (S), base (B), and large (L). For the text encoders, ${E}_{T}$ and $\bar{E}_{T}$, we use 12 layers, 512 embedding dimensions, and 8 heads, following CLIP (Radford et al., 2021). The number of text tokens is fixed to 77 with necessary truncations or paddings. Moreover, following He et al. (2021), we define our vision decoder $D_{V}$ as an 8-layer, 512-embedding dimension, and 16-head ViT. Finally, our text decoder $D_{T}$, is made of up 4 layers, 512 embedding dimension, and 8 heads.
Pre-training. We pre-train a ViT-S, ViT-B, and ViT-L using Harmony on CC3M for 25 epochs. We use an AdamW optimizer (Loshchilov & Hutter, 2019) with an $\epsilon$ value of $1e^{-6}$ to improve training stability using mixed precision (Micikevicius et al., 2018). We set the number of global crops per given image to 2 and the local crops to 8. We use a masking ratio of 75% for the pixel prediction task and pre-train up to 16 32GB V100 GPUs. Further information about our pre-training setup can be found in Appendix A.
Downstream. To asses the generalizability of learned features using our proposed method, we evaluate our model on a variety of datasets, including ImageNet-1k (Russakovsky et al., 2015) for fine-tuning, linear-probing, and zero-shot classifications, ADE20K (Zhou et al., 2017) for semantic segmentation, MS-COCO (Lin et al., 2015) for object detection and instance segmentation, as well as many other datasets for zero-shot evaluations (see Table 2). We fine-tune for 100 epochs on ImageNet-1k and for 160k iterations for ADE20K, using UperNet (Xiao et al., 2018). For object detection and instance segmentation, we use a Cascade Mask R-CNN (Cai & Vasconcelos, 2019) and fine-tune for 12 epochs. Further details on downstream settings can be found in Appendix A.
@@ 4.2 Results
(a) Scaling ViT size. The gap in 0-shot performance between Harmony and CLIP increases as we increase model size, suggesting that Harmony scales better with respect to model size.
[…]
Figure 3: Zero-shot performance comparison between Harmony and CLIP when scaling ViT size and data. (a) With increasing model size, Harmony’s advantage grows. (b) With increasing dataset size, the performance gap narrows, though Harmony maintains an edge. Appendix B shows zero-shot results on CC12M across 8 other datasets.
@@
Table 1: Comparing Harmony against baseline SSL and previous best methods. All methods are pre-trained on CC3M for 25 epochs. We used the same pre-training settings for MAE, iBOT, CLIP, SigLIP, SILC and Harmony. To reproduce SLIP, and DetailCLIP, we used their public code base (Mu et al., 2021; Monsefi et al., 2024). We reproduced MaskCLIP and SLIC ourselves since their code was not publicly available (see Appendix C for further detail on MaskCLIP). FT: fine-tuning; LIN: linear-probing; SSEG: semantic segmentation; DET: object detection; ISEG: instance segmentation.
Method ViT INet-1K ADE20K COCO
0-Shot LIN FT SSEG DET ISEG
MAE S 14.6 70.9 26.9 28.9 25.6
iBOT S 58.2 79.5 37.0 44.9 39.0
CLIP S 17.2 57.0 76.7 35.3 41.0 35.7
SigLIP S 16.0 57.0 76.8 36.2 40.8 35.6
MaskCLIP S 17.7 56.1 77.0 34.5 39.4 34.5
DetailCLIP S 18.2 50.2 77.7 28.7 42.5 37.1
SLIP S 18.1 51.5 76.5 35.0 40.0 35.0
SILC S 20.1 62.8 78.2 38.1 44.0 38.3
CLIP B 19.0 61.4 79.8 40.7 42.5 37.0
SigLIP B 18.0 61.8 79.9 40.7 42.5 37.0
MaskCLIP B 19.1 58.6 79.5 38.2 40.2 35.1
DetailCLIP B 21.1 63.2 81.6 40.9 45.5 39.5
SLIP B 19.1 63.0 80.3 42.0 42.0 36.8
SILC B 23.3 68.1 81.6 44.0 45.9 39.8
Harmony S 20.8 64.4 79.5 37.5 45.6 39.5
B 23.3 69.6 82.4 44.5 48.1 41.6
@@
We begin our evaluation of Harmony by analyzing its downstream performance for a variety of visual tasks in different settings. We start in Table 1 by presenting our main result for classification, segmentation, and detection tasks. We show that Harmony significantly outperforms CLIP and SigLIP and outperforms iBOT (Zhou et al., 2022) and MAE (He et al., 2021) in all high-level and low-level tasks evaluated. We also show that Harmony outperforms SLIP (Mu et al., 2021), MaskCLIP (Dong et al., 2023), and DetailCLIP (Monsefi et al., 2024), which are joint methods similar to ours. We also show that Harmony is compettive with SILC (Naeem et al., 2024) on zero-shot classification, but outperforms SILC on other tasks such detection, linear and fine-tuning classification. Interestingly, MAE performs substantially worse than the other methods, which could be due to the distribution of the web-scraped data in CC3M. Even then, using the pixel reconstruction objective in MAE still complements the features for Harmony (shown as an ablation in Table 4). Moreover, SLIP performs worse than expected on classification tasks, which could be due to the small batch size for SimCLR’s contrastive loss (Chen et al., 2020a) (768 vs 4096 in the original SLIP (Mu et al., 2021)). This issue is less prominent in methods that use self-distillation such as Harmony and SILC (Dong et al., 2023; Naeem et al., 2024) because they do not rely on negative examples in the batch in their self-supervised losses.
Scaling Harmony. In Figure 3 we present the performance of Harmony and CLIP as we scale model sizes from small to base and large, and as we scale data size from CC3M to CC12M. As we scale the model, the performance gap increases at an increasing rate, which is a sign that Harmony scales better than CLIP with respect to ViT size. The same trend is also shown in Figure 1, going from ViT-S to ViT-B for other tasks.
When scaling the dataset size, the performance gap between Harmony and CLIP narrows. This reduction is expected, as improvements become harder to achieve at higher accuracy levels. However, it may also suggest that CLIP benefits more from larger datasets and begins to close the gap with Harmony as the data scale increases. Appendix B shows zero-shot results on CC12M across 8 other datasets.
@@
Zero-shot evaluations across datasets. Moreover, we compare the 0-shot capabilities of Harmony to CLIP, SLIP, SILC, MaskCLIP, and DetailCLIP across diverse datasets. We focus on general and natural datasets (as opposed to domain specific like StanfordCars (Yang et al., 2015) and FGVC-Aircraft (Maji et al., 2013)), and ImageNet robustness datasets like ImageNet-R. We present our results in Table 2. Harmony outperforms all other methods in 5 out of the 8 datasets, and performs better on average compared to all methods except SILC, where it is competitive. We show qualitative examples between Harmony’s and CLIP’s zero-shot predictions in Appendix 4.
@@
Retrieval. We also evaluate Harmony for image-text zero-shot retrievals on MS-COCO (Lin et al., 2015) and Flickr30K (Young et al., 2014). The results are presented in Table 3. We observe that for top 5 (R@5) and top 10 retrievals, the gap between Harmony and the other methods grows rapidly, indicating that our method is more robust, since it still generates higher probabilities for the correct retrieval, even if it can’t retrieve it within the first examples.
@@
Table 2: Zero-shot evaluations. We compare ViT-B Harmony against CLIP, SLIP, SILC, MaskCLIP, and DetailCLIP on general and natural image benchmarks for 0-shot classification. Harmony outperforms the other methods on average.
  Average Cal101111Li et al. (2022) STL-10222Coates et al. (2011) Kin700333Carreira et al. (2022) MNIST [^4] Lecun et al. (1998) CLEVR [^5] Johnson et al. (2016) INet-A [^6] Hendrycks et al. (2021b) INet-O [^7] Hendrycks et al. (2021b) INet-R [^8] Hendrycks et al. (2021a)
CLIP 28.9 53.9 87.2 17.4 10.1 12.8 4.5 24.2 21.4
MaskCLIP 28.2 49.8 84.9 15.7 12.1 14.2 4.0 24.8 20.0
DetailCLIP 29.9 55.2 89.1 17.1 11.6 11.6 4.4 27.3 22.9
SLIP 28.3 50.5 85.2 16.2 10.8 13.3 4.9 23.7 22.2
SILC 32.7 58.6 91.5 19.8 15.1 12.4 6.9 29.2 28.4
Harmony 32.9 60.3 94.3 22.5 10.1 12.3 7.6 26.1 29.8